if the moon 202 x 10 to the 5th power miles away and the rocket travels 5.5 x 10 to the 3rd power mph, how long would this trip take?
Answer by checkley71(8403)

202*10^5/5.5*10^3=36.727*10^2=3,672.7 hours to get to the moon.